A fashion magazine claims that females like to “shop” more than males. In interviews with 1000
randomly selected females, the magazine found that 700 like to “shop”. Also, it was found that 520 out of 800
randomly selected males like to “shop”. Test the magazines claim by a 0.02 level of significance. State the
practical conclusion that refers to the information in this question.

Answer by stanbon(75887)

A fashion magazine claims that females like to “shop” more than males. In interviews with 1000 randomly selected females, the magazine found that 700 like to “shop”.
--------------------
Also, it was found that 520 out of 800 randomly selected males like to “shop”.
--------------------
Test the magazines claim by a 0.02 level of significance.
-----
Ho: p(f)-p(m) <= 0
Ha: p(f)-p(m) > 0 (claim)
------
t[(700/1000)-(520/800)] = (0.7-0.65)/sqrt[[(0.7*0.3/1000) + (0.65*0.35/800)]
= 2.2488
-------
P-value = P(t > 2.2488 when df = 1800-2)) = tcdf(2.2488,100,1798) = 0.0159
--------------
Since the p-value is less than 2%, reject Ho.
The test results support the claim.
--------------------
State the practical conclusion that refers to the information in this question.
Women like to shop more than men.
Cheers,
Stan H.
---------------